OpenStudy (anonymous):

Simplify: (x + 9) (x + 8). A. x2 + 17x + 72 B. x2 + 72x + 17 C. x2 + 17 D. x2 + 72

OpenStudy (anonymous):

\[\begin{align} \text{F.O.I.L. = First Outside Inside Last}\\ \hline \\ &(ax+b)(cx+d)\\ \ \\ \text{First:}&(\color{blue}{ax}+b)(\color{blue}{cx}+d)\to \color{blue}{ax}\cdot\color{blue}{cx}=\color{blue}{acx^2}\\ \text{Outside:}&(\color{green}{ax}+b)(cx+\color{green}{d})\to \color{green}{ax}\cdot\color{green}{d}=\color{green}{adx}\\ \text{Inside:}&(ax+\color{red}{b})(\color{red}{cx}+d)\to \color{red}{b}\cdot\color{red}{cx}=\color{red}{bcx}\\ \text{Last:}&(ax+\color{orange}{b})(cx+\color{orange}{d})\to \color{orange}{b}\cdot\color{orange}{d}=\color{orange}{bd}\\ \hline \\ (ax+b)(cx+d)&=\color{blue}{acx^2}+\color{green}{adx}+\color{red}{bcx}+\color{orange}{bd} \end{align}\]

OpenStudy (anonymous):

x2+17x+72. A. Try to do it yourself. here try this one: (x+6)(x+10)

OpenStudy (anonymous):

Use formula (x+a)(x+b) = x^2+ax +bx + ab And solve.